# EnrichEvent: Enriching Social Data with Contextual Information for Emerging Event Extraction

### v4

@@ Abstract
1 Introduction
@@
Traditionally, an event is defined as an important happening in a specific time and place [3]. With the emergence of social media, this definition expands to a happening in society that attracts a sudden burst of attention in social media, and similar to the previous explanation, it occurs in a specific time and place [6]. Intuitively, when an event occurs, quick and accurate event identification is one of the most crucial concerns for the government, aid agencies, related organizations, and even for journalists and experts [34, 17]. For instance, when a natural disaster occurs, relief organizations consider event detection as one of the ways to capture different types of information, such as the severity of the incident, the number of victims, and how to provide aid for them during natural disasters. These related organizations are keen on being informed of an event occurrence as soon as possible to take appropriate actions regarding the event [27, 28, 35]. Besides, awareness of event occurrence is vital for other purposes such as marketing, commercial, and even financial markets [25, 33]. For instance, leading commercial brands always want to discover what is happening all over the market, and this awareness, accompanying the right decisions, puts them on the path to being ahead of their rivals.
@@
Despite its value and significance, quick and accurate identification of unspecified social events has yet to be fully investigated due to the following challenges. 1) Streaming and Evolutionary. Social media services provide a streaming source of information in which people freely create content around real-world events. While working on streaming sources, processing the incoming data in a single stage is a challenging and opening problem. In addition, turning a simple happening into a hot event, which everyone will be talking about, is a complex and evolutionary phenomenon comprised of three major steps. a) An event happens in the real world. b) Observers, journalists, and ordinary people start a discussion about the occurred event on social platforms. c) The event turns into a hot event for a while or fades out at the beginning [9]. How to model the evolving characteristics of an event is an arduous task. 2) Variation in Language Expressions and Aspects of Opinions. Twitter users produce and consume information in a very informal and irregular way, and the published tweets usually include idioms, abbreviations, misspellings, irregular language, and emojis. Due to the huge and diverse user base, Twitter usually contains different words and word sequences describing the same idea. Also, in most languages, many words have various meanings in different contexts. For example, the word "fair" in a different context has multiple meanings of "carnival", "treating someone right", and "having light skin/hair". This issue makes it hard for machine learning models to understand users’ purpose. Besides, while talking about an event, users may have different opinions and express them differently. Detecting and aggregating various opinions and aspects of events is another challenge.
@@
Retrospective studies in social media and event detection have proposed practical approaches for identifying events from social data by leveraging content and network structure. For instance, frequent keyword extraction has been used to detect bursty keywords corresponding to an event, and hashtags have been leveraged to capture semantics from tweets to address the short context problem [16, 39]. Similarly, clustering methods have been employed for mitigating short and sparse contexts in social posts [1]. Further, modular decomposition and community detection have been exploited to utilize links and network structures. While much work has been devoted to event detection, more research has yet to tackle the problem of unspecified event detection and its evolution. In this paper, we design an end-to-end framework, namely EnrichEvent, to address the challenges mentioned in the previous paragraph. We define an event as a single chain of clusters where each cluster contains closely related entities both in time and semantic dimensions. In addition, we leverage contextual knowledge and structural aspects of the tweets to enrich their representation and achieve better results. Exploiting the contextual knowledge assists the model in comprehending the existing relation among the tweets and generating their representation with higher quality. We conducted several experiments to evaluate all the components of our proposed framework. The empirical results validate that EnrichEvent achieves better than various baselines based on the evaluation metrics. In summary, our main contributions are as follows:
@@
2 Related Works
@@
2.1 Identification of unspecified events
@@
2.2 Identification of predetermined events
@@
2.3 Identification of sub-events
In this category, researchers consider a specific event, e.g., a football match, and aim to identify the relevant sub-events. Designing a decent approach to find the correlations between the tweets and a given event is the main challenge of this category.  Bekoulis et al. [5] investigated the temporal aspects of the tweets and proposed a two-stage framework that detects the sub-events by analyzing their evolution over time. Moreover,  Lu et al. [20] focused on semantics to generate a better representation for tweets. In addition, they used hashtags and their n-grams to capture the correlations suitably. Though much research has been done in this category, researchers can enhance the existing frameworks differently.
Figure 1: Overview of the proposed framework: EnrichEvent. Message blocks are passed into the framework in batches corresponding to different time windows, where the first three components are processing incoming data iteratively.
3 Problem Statement
@@
Definition 2.1. Social stream $\mathcal{S}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{i-1},\mathcal{M}_{i},\ldots\}$ defined as a consecutive sequence of message blocks. In this definition, we denote message block $\mathcal{M}_{i}$ as $\mathcal{M}_{i}=\{m_{j}|1\leq j\leq|\mathcal{M}_{i}|\}$, where $|\mathcal{M}_{i}|$ is total numbers of the tweets in time period $[t_{i},t_{i_{1}})$ and $m_{j}$ denotes a single tweet.
@@
Definition 2.3. An event $e=\{c_{ij}|1\leq i\leq|\mathcal{M}_{i}|,1\leq j\leq|\mathcal{C}_{i}|\}$ is formally defined as a continuous sequence of contextually related clusters where all the clusters refer to a single event. Note that we assume each social message belongs to, at most, one event.
@@
$F:\mathcal{S}\rightarrow E=\{e_{1},e_{2},\ldots\},$ (1)
4 Methodology
@@ 4.1 Framework
4.2 Trending Data Extraction
@@
Trend detection dataset: To train the trend detection model, we started preparing an offline trend detection dataset using TwitterAPI. We have collected 1.6 million Persian tweets in different categories and tried to label them manually based on key phrases. The generated dataset has informative columns such as timestamp, tweet id, text, username, reply count, like count, and retweet count. Eventually, we utilized it for training the trend detection model. We have provided more details about the trend detection dataset in Appendix A.
@@
$\arg\min_{\mathbf{W}}\mathcal{L}(\mathbf{X},\mathbf{W},\mathbf{Y}),$ (2)
@@
4.3 Contextual Knowledge Enhancement
Working on lexical and structural aspects of the tweets is a conventional way to design event detection models. However, training a well-performing model is arduous because of the challenging characteristics of the tweets, i.e., variation in language expressions and aspects of opinions. When users debate about an occurring event on social platforms, they discuss different aspects and have diverse opinions. Therefore, we have to consider various contexts where users post tweets and aim to generate a high-quality representation for them.
In this work, we considered entities within a tweet as its representative, and concentrated on two main types of entities: Named Entities and Hashtags. Named entities are critical indicators for detecting events since an event happens at a specific time and location and includes various actors. As such, named entities demonstrate one of these attributes about an event. Hashtags, on the other hand, were incorporated into social posts intrinsically to reveal the correlation of the tweets to an event or a topic, making them invaluable indicators of tweets related to an event. For instance, the following text from our dataset contains both named entities and hashtags. The named entities indicate that the attributes of the event are Silicon Valley, USA, and 2008. In addition, #bankruptcy demonstrates that the event corresponds to bankruptcy.
@@ 4.4 Event Clustering
4.5 Event Chain Formation
@@
4.6 Event Summarization
In this stage, we try to generate summaries of the detected events by adapting the OpinionDigest [36] framework to the task of event summarization. To do so, we design a 3-step procedure. (1) Assume $t$ is a random tweet, and we could extract key phrases from tweet $t$ by considering a context window with a constant length of $\gamma$ around the appeared entities. (2) In this stage, we require an NLG model that takes a set of key phrases as input and produces a fluent summary as output. Because we do not access the gold-standard summaries for training, we attempt to fine-tune a language model that encodes the extracted key phrases of a single tweet and then try to reconstruct the tweet from them. To do so, we build a synthetic dataset by mapping the existing tweets in our event detection dataset to their corresponding key phrases, similar to the first step. Furthermore, we fine-tune the GPT-2 language model on the synthetic dataset. Now, we can utilize the fine-tuned model for generating event summaries. (3) Suppose that we want to generate a summary for event $e$, and $T=\{t_{1},t_{2},\ldots,t_{n}\}$ is set of corresponding tweets to event $e$. First, we map all the tweets to their corresponding key phrases. Then, we leverage the DBSCAN model to cluster the key phrases using their BERT representation. In each cluster, we select the closest key phrase to the center of the clusters. Eventually, we utilize the fine-tuned GPT-2 model to generate the summary of event $e$ using the selected key phrases.
4.7 Storage and Evaluation
After generating the summaries, we store the details of the identified events as JSON files in the database. While saving the results, we extract informative details of the events, such as event ID, event period, event summary, details of corresponding entities/hashtags/usernames to each event, and tweets with the highest like/reply/retweet count. Concisely, our proposed framework gets the message blocks as input and returns an informative JSON file for each detected event. In the final stage, we evaluate the performance of our proposed framework from various aspects. We discuss the results and evaluation details in the following section.
5 Experiments and Evaluation Pipeline
@@
5.1 Experimental Setup
5.1.1 Dataset
Similar to the trend detection dataset, we have prepared a dataset for evaluating the overall performance of EnrichEvent on Twitter streaming data using TwitterAPI. We considered a two-week time window and crawled the posted Persian tweets from 2023/2/28 to 2023/3/14. We collected almost 760000 tweets in this period and divided them into daily message blocks. Since we had to label the entities manually, we sampled 500 tweets in each message block. We selected the tweets based on the words’ frequency to preserve the actual distribution of tweets. In other words, if a tweet contains frequent words, the probability of choosing it will be higher. We provided more details about the event detection dataset in Appendix A.
5.1.2 Baselines
@@ 5.1.3 Evaluation Metrics
$\begin{gathered}\mathcal{C}(e)=\frac{1}{|T_{e}|}\sum_{t\in T_{e}}\frac{\mathcal{A}_{t}}{\mathcal{U}_{t}}\end{gathered}$ (3)
@@
$\begin{gathered}\mathcal{D}(e)=\frac{1}{|T_{e}|}\sum_{t\in T_{e}}\frac{\mathcal{B}_{t}}{\mathcal{U}_{t}}\end{gathered}$ (4)
5.1.4 Implementation Details
We have implemented our proposed framework, i.e., EnrichEvent, using a machine equipped with a Tesla T4 GPU, and its source code is publicly available on Github. Details of the hyper-parameters are shown in Table 1.
Table 1: Details of the hyper-parameters
Hyper-parameter
Description
Value
$\sigma$
Minimum threshold for the probability of referring to an event in the Trend Detection component
$0.3$
$\alpha$
Weight of the entity-to-entity distance matrix obtained from the Embedding matrix
$0.5$
$\beta$
Total number of selected tweets for generating the embedding vectors for entities
$4$
$\gamma$
length of the content window for extracting key phrases in the Event Summarization component
$2$
$\delta$
Minimum number of common entities to add an edge among clusters in the Event Chain Formation component
$3$
5.2 Manual Evaluation
We evaluated the overall performance of EnrichEvent and the selected baselines on Twitter streaming message blocks and presented the results in Table 2. To calculate the evaluation metrics, we manually labeled the entities. Note that Consolidation and Discrimination are calculated for a single event, and the average results obtained from all the detected events are presented as the final result.
Table 2: Manual evaluation on streaming social data
Method Consolidation Discrimination
@@ 5.2 Manual Evaluation
Discussion: As shown in Table 2, the results indicate that our proposed framework outperforms all the baselines. We should mention some key findings from the results are as follows: (1) In section 4.3, we claimed that if we leverage the contextual and lexical knowledge simultaneously, it will empower the event detection frameworks. The obtained results validate the superiority of our proposed framework compared to other baselines. (2) Results of Yang et al. reveals that hashtags are not good representatives of tweets. They do not transfer considerable knowledge for event detection since users do not usually use hashtags while posting tweets. (3) Comparing the results of Fedoryszak et al. and BERT demonstrate that exploiting the contextual knowledge reduces the discrimination metric surprisingly by 10%, but it does not affect the value of the Consolidation individually. (4) Comparing the results of BERT and EnrichEvent reveals that utilizing contextual knowledge along with the lexical aspects increases the ratio of related entity pairs in cluster chains by 20.30% and is the key to achieving the best results.
5.3 Clusters’ Quality
@@
As shown in Figure 2, EnrichEvent yields the highest fraction of related entities. That is because it fully uses contextual and lexical representations of the tweets. Compared to Fedoryszak et al., which only utilized co-occurrence vectors as their representation, we achieve 13.98% promotion in the fraction of related entities. It conveys that EnrichEvent gets a good comprehension of the events and clusters them properly. To validate the quality of clusters, we also investigate the clustering metrics.
@@
5.4 Evaluation of Trending Data Extraction

### v5

@@ Abstract
Index Terms:
Unspecified Event Detection, Dynamic Event Detection, Computational Social Science, Social Networks, Twitter
I Introduction
@@
Traditionally, an event is defined as an important happening in a specific time and place [5]. With the emergence of social media, this definition expands to a happening in society that attracts a sudden burst of attention in social media, and similar to the previous explanation, it occurs in a specific time and place [6]. Intuitively, when an event occurs, quick and accurate event identification is one of the most crucial concerns for the government, aid agencies, related organizations, and even for journalists and experts [7, 8]. For instance, when a natural disaster occurs, relief organizations consider event detection as one of the ways to capture different types of information, such as the severity of the incident, the number of victims, and how to provide aid for them during natural disasters. These related organizations are keen on being informed of an event as soon as possible to take appropriate actions regarding the event [9, 10, 11]. Besides, awareness of event occurrence is vital for other purposes such as marketing, commercial, and even financial markets [12, 13]. For instance, leading retail brands always want to discover what is happening all over the market, and this awareness, accompanying the right decisions, puts them on the path to being ahead of their rivals.
@@
Despite its value and significance, quick and accurate identification of unspecified social events has yet to be fully investigated due to the following challenges. 1) Streaming and Evolutionary. Social media services provide a streaming source of information in which people freely create content around real-world events. While working on streaming sources, processing the incoming data in a single stage is a challenging and opening problem. In addition, turning a simple happening into a hot event, which everyone will be talking about, is a complex and evolutionary phenomenon comprised of three major steps. a) An event happens in the real world. b) Observers, journalists, and ordinary people start a discussion about the occurred event on social platforms. c) The event turns into a hot event for a while or fades out at the beginning [1]. How to model the evolving characteristics of an event is an arduous task. 2) Variation in Language Expressions and Aspects of Opinions. Twitter users produce and consume information in a very informal and irregular way, and the published tweets usually include idioms, abbreviations, misspellings, irregular language, and emojis. Due to the huge and diverse user base, Twitter usually contains different words and word sequences describing the same idea. Also, in most languages, many words have various meanings in different contexts. For example, the word ”fair” in a different context has multiple meanings, such as ”carnival,” ”treating someone right,” and ”having light skin/hair.” This issue makes it hard for machine learning models to understand users’ purpose. Besides, while talking about an event, users may have different opinions and express them differently. Detecting and aggregating various opinions and aspects of events is another challenge.
@@
Retrospective studies in social media and event detection have proposed practical approaches for identifying events from social data by leveraging content and network structure. For instance, frequent keyword extraction has been used to detect bursty keywords corresponding to an event, and hashtags have been leveraged to capture semantics from tweets to address the short context problem [14, 15]. Similarly, clustering methods have been employed for mitigating short and sparse contexts in social posts [16]. Further, modular decomposition and community detection have been exploited to utilize links and network structures. While much work has been devoted to event detection, more research has yet to tackle the problem of unspecified event detection and its evolution. In this paper, we design an end-to-end framework, namely EnrichEvent, to address the challenges mentioned in the previous paragraph. We define an event as a single chain of clusters where each cluster contains closely related entities both in time and semantic dimensions. In addition, we leverage contextual knowledge and structural aspects of the tweets to enrich their representation and achieve better results. Exploiting contextual knowledge assists the model in comprehending the existing relation among tweets and generating a higher-quality representation. We conducted several experiments to evaluate all the components of our proposed framework. The empirical results validate that EnrichEvent achieves better than various baselines based on the evaluation metrics. In summary, our main contributions are as follows:
@@
Figure 1: Overview of the proposed framework: EnrichEvent. Message blocks are passed into the framework in batches corresponding to different time windows, where we first filter out the pointless tweets and extract the existing named entities within the tweets. Then, we categorize the extracted named entities into clusters based on their lexical and semantical representation. Finally, we connect the related clusters in consecutive message blocks and represent each event as a cluster chain. As an output, EnrichEvent stores the information of each event in a JSON file.
II Related Works
@@
II-A Identification of unspecified events
@@
II-B Identification of predetermined events
@@
II-C Identification of sub-events
In this category, researchers consider a specific event, e.g., a football match, and aim to identify the relevant sub-events. Designing a decent approach to find the correlations between the tweets and a given event is the main challenge of this category.  Bekoulis et al. [28] investigated the temporal aspects of the tweets and proposed a two-stage framework that detects the sub-events by analyzing their evolution over time. Moreover,  Lu et al. [29] focused on semantics to generate a better representation of tweets. In addition, they used hashtags and their n-grams to capture the correlations suitably. Though much research has been done in this category, researchers can enhance the existing frameworks differently.
III Problem Statement
@@
Definition 2.1. Social stream $\mathcal{S}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{i-1},\mathcal{M}_{i},\ldots\}$ defined as a consecutive sequence of message blocks. We denote message block $\mathcal{M}_{i}$ as $\mathcal{M}_{i}=\{m_{j}|1\leq j\leq|\mathcal{M}_{i}|\}$, where $|\mathcal{M}_{i}|$ is total numbers of the tweets in time period $[t_{i},t_{i_{1}})$ and $m_{j}$ denotes a single tweet.
@@
Definition 2.3. $e=\{c_{ij}|1\leq i\leq|\mathcal{M}_{i}|,1\leq j\leq|\mathcal{C}_{i}|\}$ is a social event and formally defined as a continuous sequence of contextually related clusters where all the clusters refer to a single event. Note that we assume each social message belongs to, at most, one event.
@@
$\centering F:\mathcal{S}\rightarrow E=\{e_{1},e_{2},\ldots\}\@add@centering$ (1)
IV Methodology
@@ IV-A Framework
IV-B Trending Data Extraction
@@
Trend detection dataset: To train the trend detection model, we started preparing an offline trend detection dataset using TwitterAPI. We have collected 1.6 million Persian tweets in different categories and tried to label them manually based on key phrases. The generated dataset has informative columns such as timestamp, tweet id, text, username, reply count, like count, and retweet count. Eventually, we utilized it to train the trend detection model. We have provided more details about the trend detection dataset in Appendix A.
@@
$\centering\arg\min_{\mathbf{W}}\mathcal{L}(\mathbf{X},\mathbf{W},\mathbf{Y}),\@add@centering$ (2)
@@
IV-C Contextual Knowledge Enhancement
Working on lexical and structural aspects of tweets is a conventional way to design event detection models. However, training a well-performing model is arduous because of the tweets’ challenging characteristics, i.e., variation in language expressions and aspects of opinions. When users debate about an occurring event on social platforms, they discuss different aspects and have diverse opinions. Therefore, we have to consider various contexts where users post tweets and aim to generate a high-quality representation for them.
In this work, we considered entities within a tweet as its representative and concentrated on two main types of entities: Named Entities and Hashtags. Named entities are critical indicators for detecting events since an event happens at a specific time and location and includes various actors. As such, named entities demonstrate one of these attributes about an event. Hashtags, on the other hand, were intrinsically incorporated into social posts to reveal the correlation of the tweets to an event or a topic, making them invaluable indicators of tweets related to an event. For instance, the following text from our dataset contains both named entities and hashtags. The named entities indicate that the attributes of the event are Silicon Valley, USA, and 2008. In addition, #bankruptcy demonstrates that the event corresponds to bankruptcy.
@@ IV-D Event Clustering
IV-E Event Chain Formation
@@
IV-F Event Summarization
In this stage, we try to generate summaries of the detected events by adapting the OpinionDigest [34] framework to the task of event summarization. To do so, we design a 3-step procedure. (1) Assume $t$ is a random tweet, and we could extract key phrases from tweet $t$ by considering a context window with a constant length of $\gamma$ around the appeared entities. (2) In this stage, we require an NLG model that takes a set of key phrases as input and produces a fluent summary as output. Because we do not access the gold-standard summaries for training, we attempt to fine-tune a language model that encodes the extracted key phrases of a single tweet and then try to reconstruct the tweet from them. To do so, we build a synthetic dataset by mapping the existing tweets in our event detection dataset to their corresponding key phrases, similar to the first step. Furthermore, we fine-tune the GPT-2 language model on the synthetic dataset. Now, we can utilize the fine-tuned model to generate event summaries. (3) Suppose that we want to generate a summary for event $e$, and $T=\{t_{1},t_{2},\ldots,t_{n}\}$ is set of corresponding tweets to event $e$. First, we map all the tweets to their corresponding key phrases. Then, we leverage the DBSCAN model to cluster the key phrases using their BERT representation. In each cluster, we select the closest key phrase to the center of the clusters. Eventually, we utilize the fine-tuned GPT-2 model to generate the summary of event $e$ using the selected key phrases.
IV-G Storage and Evaluation
After generating the summaries, we store the details of the identified events as JSON files in the database. While saving the results, we extract informative information about the events, such as event ID, event period, event summary, details of corresponding entities/hashtags/usernames to each event, and tweets with the highest like/reply/retweet count. Concisely, our proposed framework gets the message blocks as input and returns an informative JSON file for each detected event. In the final stage, we evaluate the performance of our proposed framework from various aspects. We discuss the results and evaluation details in the following section.
V Experiments and Evaluation Pipeline
@@
V-A Experimental Setup
V-A1 Dataset
Similar to the trend detection dataset, we have prepared a dataset to evaluate the overall performance of EnrichEvent on Twitter streaming data using TwitterAPI. We considered a two-week time window and crawled the posted Persian tweets from 2023/2/28 to 2023/3/14. We collected almost 760000 tweets in this period and divided them into daily message blocks. Since we had to label the entities manually, we sampled 500 tweets in each message block. We selected the tweets based on the words’ frequency to preserve the actual distribution of tweets. In other words, if a tweet contains frequent words, the probability of choosing it will be higher. We provided more details about the event detection dataset in Appendix A.
V-A2 Baselines
@@ V-A3 Evaluation Metrics
$\centering\mathcal{C}(e)=\frac{1}{|T_{e}|}\sum_{t\in T_{e}}\frac{\mathcal{A}_{t}}{\mathcal{U}_{t}}\@add@centering$ (3)
@@
$\centering\mathcal{D}(e)=\frac{1}{|T_{e}|}\sum_{t\in T_{e}}\frac{\mathcal{B}_{t}}{\mathcal{U}_{t}}\@add@centering$ (4)
V-A4 Implementation Details
We have implemented our proposed framework, EnrichEvent, using a machine equipped with a Tesla T4 GPU. The replication package for the EnrichEvent framework is publicly available on Github [37], and you can use it to reproduce the results. Details of the hyper-parameters are shown in Table I.
TABLE I: Details of the hyper-parameters
Hyper-parameter Description Value
$\sigma$ Minimum threshold for the probability of referring to an event in the Trend Detection component $0.3$
$\alpha$ Weight of the entity-to-entity distance matrix obtained from the Embedding matrix $0.5$
$\beta$ Total number of selected tweets for generating the embedding vectors for entities $4$
$\gamma$ length of the content window for extracting key phrases in the Event Summarization component $2$
$\delta$ Minimum number of common entities to add an edge among clusters in the Event Chain Formation component $3$
V-B Manual Evaluation
We evaluated EnrichEvent’s overall performance and the selected baselines on Twitter streaming message blocks and presented the results in Table II. To calculate the evaluation metrics, we manually labeled the entities. Note that Consolidation and Discrimination are calculated for a single event, and the average results obtained from all the detected events are presented as the final result.
TABLE II: Manual evaluation on streaming social data
[…]
@@ V-B Manual Evaluation
Discussion: As shown in Table II, the results indicate that our proposed framework outperforms all the baselines. We should mention some key findings from the results are as follows: (1) In section 4.3, we claimed that if we leverage the contextual and lexical knowledge simultaneously, it will empower the event detection frameworks. The obtained results validate the superiority of our proposed framework compared to other baselines. (2) Results of Yang et al. reveals that hashtags are not good representatives of tweets. They do not transfer considerable knowledge for event detection since users do not usually use hashtags while posting tweets. (3) Comparing the results of Fedoryszak et al. and BERT demonstrate that exploiting the contextual knowledge reduces the discrimination metric surprisingly by 10%, but it does not affect the value of the Consolidation individually. (4) Comparing the results of BERT and EnrichEvent reveals that utilizing contextual knowledge along with the lexical aspects increases the ratio of related entity pairs in cluster chains by 20.30% and is the key to achieving the best results.
V-C Clusters’ Quality
@@
As shown in Figure 2, EnrichEvent yields the highest fraction of related entities compared to the baselines. That is because it fully uses contextual and lexical representations of the tweets. Compared to Fedoryszak et al., which only utilized co-occurrence vectors of named entites as their representation, we achieve 13.98% promotion in the fraction of related entities. It conveys that EnrichEvent achieved a good comprehension of the events and clusters them properly. To validate the quality of clusters, we also investigated the clustering metrics which measure the quality of clusters in different perspective.
@@
V-D Evaluation of Trending Data Extraction